# Metal-Poor Stars in the MW Disk: Resonant Cooling of Vertical Oscillations of Halo Stars in Barred Galaxies

### v1

@@ 1 Introduction
Recent observations have detected a population of low metallicity stars with [Fe/H]$\leq-2.5$ dex in the Milky Way (MW) stellar disk. About 31% of these stars reside in the galactic disk, within $|z|\leq 3$ kpc of its midplane (e.g., Sestito et al., 2020; Lucey et al., 2021). The Pristine photometric survey (Starkenburg et al., 2017), Hamburg/ESO (Reimers, 1990; Beers et al., 1999) and Gaia EDR3 (Gaia Collaboration et al., 2021) surveys have quantified some of the properties of this population. Moreover, this population appears to be statistically biased to populate prograde orbits with respect to the disk rotation. Bellazzini et al. (2023) have analyzed the metal-poor stars in the Solar neighborhood and found that among the disky kinematics, the ratio of the prograde to retrograde orbits is $\sim 3/1$. Finally, Zhang et al. (2023) have analyzed the Gaia DR3 kinematic sample of the metal-poor population in the MW and found two kinematic groups — one stationary and one with a net prograde rotation of $80\,{\rm km\,s^{-1}}$.
@@
Several suggestions have been made to the origin of these metal-poor stars in the disk. It can come from the early buildup of the MW, especially stars on the retrograde orbits, and from minor mergers (e.g., Sestito et al., 2021; Carollo et al., 2023). Metal-poor disk stars also appear on very eccentric orbits. Other suggestions include the early formation of the metal-poor stars entirely in the MW disk that is heated to form a thick disk (e.g., Di Matteo et al., 2020). Or they could be shepherded by the stellar bar, i.e., trapped by its corotation resonance from the inner galaxy (e.g., Yuan et al., 2023).
@@
Here, we analyze a new process which affects the halo population, moving a small fraction of this population into the stellar disk, and propose a novel scenario for the origin of the observed low metallicity stellar population confined close to the mid-plane of the MW and other disk galaxies. The majority of disk galaxies host stellar bars on various spatial scales (e.g., Sellwood & Wilkinson, 1993; Knapen et al., 2000; Jogee et al., 2004). These systems have been modeled extensively using 3-D numerical simulations, and many of their properties have been successfully analyzed (e.g., Sellwood, 1980; Athanassoula et al., 1983; Pfenniger, 1984; Weinberg, 1985; Combes et al., 1990; Raha et al., 1991; Athanassoula, 2003; Martinez-Valpuesta et al., 2006; Dubinski et al., 2009; Villa-Vargas et al., 2009; Bi et al., 2022). Nevertheless, these collisionless systems continue to provide new effects, such the effect of spinning DM halos and bulges (e.g., Long et al., 2014; Collier et al., 2018; Li et al., 2023b, c). We define the DM halo spin as $\lambda=J/J_{\rm max}$ following Bullock et al. (2001), where $J$ is the halo angular momentum and $J_{\rm max}$ is maximally allowed $J$.
Interactions between the stellar population of galactic disks and galactic bars have been studied extensively and are dominated by resonances. These interactions include resonances between the orbits in the disk plane and those in the vertical plane have been found to be associated with the energy transfer from the disk motions to the vertical motions (Friedli & Pfenniger, 1990; Combes et al., 1990; Li et al., 2023a). More generally, one expects that interactions between the DM halo population and the embedded barred stellar disks involve resonances between halo orbits and stellar bars. In this work, we highlight one of such resonant effects which leads to the cooling of vertical stellar oscillations in spinless and spinning halos. We limit ourselves to the outcomes of this effect — the origin of disk metal-poor stars. In the follow-up work, we will take a more comprehensive look at halo-disk resonances.
Collisionless systems, such as galaxies or DM halos exchange their energies on a timescale which is much longer than the Hubble time. Their relaxation in this process characterized by small angle encounters, i.e., essentially two-body distant collisions, can be described using the Fokker-Planck equation (e.g., Binney & Tremaine, 2008). The alternative to this relaxation is the evolution (i) in a time-dependent gravitational potential, when the individual energies of particles are not conserved and can be exchanged with the mean field (Lynden-Bell, 1967), or (ii) by resonant interactions. The former evolution does not lead to the Maxwellian distribution of velocities, while the end-process of the latter interaction is less clear.
In this work, we use two standard models, with $\lambda=0$ and 0.06, in order to demonstrate an interesting effect, which so far has eluded the attention of modelers and theorists. The higher value of $\lambda$ has been selected based on the first measurement of the MW DM halo spin, $\lambda\sim 0.061-0.088$ (Obreja et al., 2022). Section 2 describes the numerical models, section 3 presents our results, and section 4 analyzes these results and discusses their implications for galaxy evolution.
@@ 2 Numerical Methods
We use the $N$-body part of the mesh-free hydrodynamics code gizmo (Hopkins, 2015), an extension of the gadget-2 code (Springel, 2005). The models of stellar disk galaxies embedded in spherical DM halos have been constructed with two different halo spins using methods described elsewhere (e.g., Collier et al., 2018; Li et al., 2023b). We re-simulated the model with the halo virial radius of 180 kpc from Li et al. (2023b) with $\lambda=0$ and $\lambda=0.06$, using $N_{\mathrm{DM}}=7.2\times 10^{6}$ DM particles and $N_{\mathrm{S}}=8\times 10^{5}$ stellar particles. The mass-per-particle in the disk is $7.9\times 10^{4}\,\mathrm{M_{\odot}}$ and $8.7\times 10^{4}\,\mathrm{M_{\odot}}$ in the halo. The gravitational softening is $\epsilon=25\,\mathrm{pc}$ for each particle. The models have been run for $10\,\mathrm{Gyr}$, with total angular momentum conserved within 0.2% and total energy within 0.1% over the total integration time.
The DM component has the NFW density profile (Navarro et al., 1996) relaxed in the exponential disk potential, with $r_{\mathrm{c}}=1.4\,\mathrm{kpc}$, the flat density core, and $r_{\mathrm{s}}=10\,\mathrm{kpc}$, the characteristic radius. We use a Gaussian cut-off radius $r_{\mathrm{t}}=180$ kpc to obtain a finite mass of $M_{\mathrm{h}}=6.3\times 10^{11}\,\mathrm{M_{\odot}}$. The initial exponential stellar disk has the mass $M_{\mathrm{d}}=6.3\times 10^{10}\,\mathrm{M_{\odot}}$ and radial scale-length $R_{0}=2.85$ kpc, and the scale-height $z_{0}=0.6\,\mathrm{kpc}$. Velocities of disk particles are assigned using the epicycle approximation and asymmetric drift correction (Li et al., 2023b).
Two models of DM halos are abbreviated as P00 (for $\lambda=0$) and P60 (for $\lambda=0.06$). The angular momentum axis of the DM halo in the P60 model agrees with that of the disk.
@@
Figure 2: A schematic figure of the region where the DM particles are sampled. The Figure is in the disk plane’s edge-on view ($x-z$ plane) at $t=0.5$ Gyr. The colored contours are the equipotential surfaces of the total system. The gray dashed box indicates the volume of a cylinder with a radius $R=15$ kpc and vertical extension from $z=-15$ to $z=15$ kpc. The dashed ‘circle’ is an equipotential surface with the maximal potential $\Phi_{\mathrm{max}}$ found inside the cylinder. The envelope of the disk is plotted as the solid black curve. We randomly sampled $6.5\times 10^{5}$ DM particles inside $\Phi_{\mathrm{max}}$ equipotential surface whose specific energy is smaller than the $\Phi_{\mathrm{max}}$.
@@ 3 Results
Both models, P00 and P60, are subject to the bar instability and their evolution is described in Li et al. (2023b) and in Figure 1, displaying the bar Fourier amplitude, $A_{2}$, the buckling Fourier amplitude, $A_{\rm 1z}$, and the bar pattern speed, $\Omega_{\rm bar}$. The stellar bar evolution depends on the halo $\lambda$. The P00 bar experiences vertical buckling at $t\sim 3.3$ Gyr, and fully re-grows after, while the P60 bar buckles at about 4 Gyr, essentially leaving behind only an oval distortion, and its torque against the DM, therefore, is very weak.
@@
We have performed an orbit analysis [see e.g., Li et al. (2023c) for details] for both models by constructing a sample of halo particles from each model. The sampled particles have been selected as follows (see Figure 2). (1) We calculated the maximal potential $\Phi_{\mathrm{max}}$ from the region within the cylindrical radius $R<15$ kpc and vertical distance $|z|<15$ kpc at $t=0.5$ Gyr; (2) inside the equipotential surface of $\Phi_{\mathrm{max}}$, we randomly choose $6.5\times 10^{5}$ DM particles whose specific energy is smaller than $\Phi_{\mathrm{max}}$.
@@
We have determined various parameters of individual orbits of DM particles, namely, their maxima along the $z$-axis, $z_{\rm max}$, and along the cylindrical $R$-axis, $R_{\rm max}$. By focusing on the evolution of $z_{\rm max}(t)$, we have detected its decrease by a factor $\geq 2$ over various time periods. To analyze the statistical properties of these DM particle orbits, we have binned the evolution of $z_{\rm max}(t)$ into time periods of $\Delta t\leq 0.5$ Gyr. We call abruptly ‘cooled’ particles those cooled over $\Delta t\leq 0.5$ Gyr, to separate them from all other cooled halo particles over longer time periods, i.e., 0.5–1.0 Gyr, 1.0–1.5 Gyr, 1.5–2.0 Gyr, and so on.
Note that the $z$-coordinate of each particle has been measured with respect to the Laplace surface defined by the surface where the vertical acceleration $a_{z}$ vanishes (e.g., Li et al., 2023c). Next, we have calculated $\langle z_{\mathrm{max}}\rangle$ —the time-averaged $z_{\mathrm{max}}$ for each consecutive 0.5 Gyr bin from $t=0$ to $t=10$ Gyr.
@@
Figure 3: Four representative examples of cooling particles in the P00 model. The $z$-coordinate of each particle is measured with respect to the Laplace surface. (a) Abruptly cooling particle at $\sim 6.5$ Gyr and never heats up again until the end of the run. (b) The particle cools down abruptly at $\sim 3.5$ Gyr and heats up at $\sim 8$ Gyr. (c) The particle experiences multiple coolings at $\sim 2.5$, $\sim 5.8$, and $\sim 8$ Gyr. (d) The particle cools down gradually from $\sim 2$ to $\sim 10$ Gyr.
@@
Figure 4: Statistics of a sub-sample of DM halo particles which cool down from high altitude into the disk plane. These particles are selected under two conditions (see text for details) from the P00 (blue) and the P60 (red) models. Top frame: The net number of sub-sampled particles whose $\langle z_{\mathrm{max}}\rangle$ are smaller than the disk thickness after a cooling event. Particles that are heated up and leaving the disk (i.e., $\langle z_{\mathrm{max}}\rangle$ larger than the disk thickness) in each time bin are excluded. Bottom frame: The histogram of the duration of individual cooling events for particles that are found in the disk at $t=10$ Gyr (i.e., the particles at the end-point in the top panel).
@@
To obtain the statistics of cooled halo particles, we proceed with the following definitions. First, we assume that the stellar disk extends to $z_{\rm disk}=3$ kpc above/below the midplane, based on the edge-on disk density contours. Second, we focus on these halo particles which acquire the disk kinematics. Third, in Section 4, we consider a fraction of cooled DM particles as the low metallicity halo stars that have obtained the disk kinematics. For this purpose, we limit our statistical analysis to DM particles with $\langle z_{\mathrm{max}}\rangle\geq 2z_{\rm disk}$ only before the cooling process begins. This also explains the requirement of cooling $\langle z_{\mathrm{max}}\rangle$ by a factor of $\geq 2$.
We define a sub-sample of all DM halo particles that cool down from the high altitude into the disk in each bin by all types of cooling, i.e., abruptly, gradually, and multiple coolings, $N_{\rm cool}(t)$. In both the P00 and P60 models, the particles of this sub-sample are filtered out of the total sample discussed above by imposing the two conditions: (1) $\langle z_{\mathrm{max}}\rangle>2z_{\rm disk}$ at the time of sampling ($t=0.5\>\mathrm{Gyr}$); (2) $\langle z_{\mathrm{max}}\rangle<z_{\rm disk}$, after the cooling.
The top panel of Figure 4 shows the cumulative number of cooled DM halo particles as a function of time, $N_{\rm cool}(t)$, for each model, i.e., only the DM particles that remain in the disk. The re-heated DM particles, i.e., those leaving the disk, have been excluded. Hence, $N_{\rm cool}(t=10\,\rm Gyr)$ gives the number of the cooled DM particles remaining in the disk by the end of the run. The bottom panel shows the duration distribution of individual cooling events over $t=10$ Gyr.
The sample of DM particles used to analyze the statistics of cooling events includes only the particles with $\langle z_{\mathrm{max}}\rangle$ larger than two times $z_{\rm disk}$. This number is $N_{\rm total}=325,270$ for the P00 model and 325,019 for the P60 model.

### v2

@@ 1 Introduction
Recent observations have detected a population of low metallicity stars with [Fe/H]$\leq$-2.5 dex in the Milky Way (MW) stellar disk. 31% of these stars reside in the disk, $|z|\leq 3$ kpc of its midplane (e.g., Sestito et al., 2020; Lucey et al., 2021). Pristine photometric survey (Starkenburg et al., 2017), Hamburg/ESO (Reimers, 1990; Beers et al., 1999) and Gaia EDR3 (Gaia Collaboration et al., 2021) surveys quantified some of the properties of this population. It appears to be statistically biased to populate prograde (with disk rotation) orbits. Bellazzini et al. (2024) have analyzed metal-poor stars in the Solar neighborhood and found that among the disky kinematics, the ratio of the prograde to retrograde orbits is $\sim 3/1$. Zhang et al. (2023) have analyzed the Gaia DR3 kinematic sample of the metal-poor population and found two kinematic groups — one stationary and one with a net prograde rotation of $80\,\mathrm{km\,s^{-1}}$.
@@
Suggestions have been made to the origin of these metal-poor stars in the disk. They can come from early buildup of the MW, especially retrograde stars, and from minor mergers (e.g., Sestito et al., 2021; Carollo et al., 2023). Metal-poor disk stars also appear on very eccentric orbits. Other suggestions include the early formation of these stars entirely in the MW disk that is heated to a thick disk (e.g., Di Matteo et al., 2020). Or they could be shepherded by the stellar bar, i.e., trapped by its corotation resonance from the inner galaxy (e.g., Dillamore et al., 2023; Yuan et al., 2023).
@@
Here, we analyze a new process which affects the halo population, moving its small fraction into the stellar disk, and propose a novel scenario for the origin of the observed low metallicity stars confined close to the mid-plane of disk galaxies. Majority of disk galaxies host stellar bars on various sizes (e.g., Sellwood & Wilkinson, 1993; Knapen et al., 2000; Jogee et al., 2004). These systems have been modeled extensively using 3D numerical simulations, and many of their properties have been successfully analyzed (e.g., Sellwood, 1980; Athanassoula et al., 1983; Pfenniger, 1984; Weinberg, 1985; Combes et al., 1990; Raha et al., 1991; Athanassoula, 2003; Martinez-Valpuesta et al., 2006; Dubinski et al., 2009; Villa-Vargas et al., 2009; Bi et al., 2022). Nevertheless, these collisionless systems continue to provide new effects, e.g., effect of spinning halos and bulges (e.g., Long et al., 2014; Collier et al., 2018; Li et al., 2023b, c). We define the DM halo spin as $\lambda=J/J_{\mathrm{max}}$ following Bullock et al. (2001), where $J$ is the halo angular momentum and $J_{\mathrm{max}}$ is maximally allowed $J$.
The 3-D dynamics of barred galaxies was found to be associated with a substantially chaotic phase space, whose fraction correlates with the bar strength (Athanassoula et al., 1983; Pfenniger, 1984) and the host halo departure from axial symmetry (El-Zant & Shlosman, 2002). In conservative dynamical systems, the size of chaotic region is a measure of the resonance width — both the concepts of a chaos and a resonance are related. Resonances can overlap, so it becomes impossible to relate a given chaotic orbit to single resonance. In more than 2-D non-integrable potentials, the chaotic space at a given energy is connected via the Arnold’s web “surrounding” the regular islands of quasi-periodic orbits (Lichtenberg & Lieberman, 1992). In a secular evolution, the regular islands may grow or shrink, allowing orbits to migrate in a stochastic way between islands possessing different orbital properties.
One can partition the phase space in barred galaxies in three distinct regions (see Fig. 1 in Ollé & Pfenniger, 1998): first, inside the corotation, at the Jacobi energy lower than the one of the relevant Lagrange points; second, outside the corotation; and third, the rest of the phase space at higher Jacobi energy. The last region involves most of the DM particles, whose motion is allowed throughout the phase space. The fraction of chaotic phase space in the bar neighborhood is expected to be large, as very few regular islands exist, e.g., the retrograde inclined “anomalous” family (Heisler et al., 1982). In particular, the radial orbits along the bar rotation axis with amplitudes exceeding the disk scaleheight are very unstable (Martinet & Pfenniger, 1987; Pfenniger, 1987), meaning that the phase space close to the rotation axis in the halo is connected to the chaotic web.
The most chaotic halo orbits are those able to oscillate between large radii and the bar region. Away from the bar, they behave as orbits in the axisymmetric or spherical potentials. When penetrating the bar region they can be scattered quasi-randomly, subject to bar orientation. Resonances have been found to dominate interactions between the vertical and disk orbits, and are associated with energy transfer away from the disk motions (Friedli & Pfenniger, 1990; Combes et al., 1990; Li et al., 2023a). Moreover, interactions between the halo population and the embedded barred stellar disks involve many resonances between halo orbits and stellar bars, connecting widely different phase space regions by the chaotic web.
In this work, we highlight one of such effects that has eluded the attention of modelers and theorists — the resonant cooling of vertical stellar oscillations in spinless and $\lambda=0.06$ halos. The higher value of $\lambda$ has been selected based on the measurement of the MW DM halo spin, $\lambda\sim 0.061-0.088$ (Obreja et al., 2022). We limit ourselves to the by-product of this effect — the origin of disk metal-poor stars. We analyze our results and discuss their implications for galaxy evolution.
@@ 2 Numerical Methods
We use the $N$-body part of the mesh-free hydrodynamics code gizmo (Hopkins, 2015), an extension of the gadget-2 code (Springel, 2005). The models of stellar disk galaxies embedded in spherical DM halos have been constructed with two different halo spins using methods described elsewhere (e.g., Collier et al., 2018; Li et al., 2023b). We re-simulated the model with the halo virial radius of 180 kpc from Li et al. (2023b) with $\lambda=0$ and $\lambda=0.06$, increasing the frequency of dumping snapshots to every 1 Myr, using $N_{\mathrm{DM}}=7.2\times 10^{6}$ DM particles and $N_{\mathrm{S}}=8\times 10^{5}$ stellar particles. The mass-per-particle in the disk is $7.9\times 10^{4}\,\mathrm{M_{\odot}}$ and $8.7\times 10^{4}\,\mathrm{M_{\odot}}$ in the halo. The gravitational softening is $\epsilon=25\,\mathrm{pc}$ for each particle. The models have been run for $10\,\mathrm{Gyr}$, with total angular momentum conserved within 0.2% and total energy within 0.1% over the total integration time.
The DM component has the NFW density profile (Navarro et al., 1996) relaxed in the exponential disk potential, with $r_{\mathrm{c}}=1.4\,\mathrm{kpc}$, the flat density core, and $r_{\mathrm{s}}=10\,\mathrm{kpc}$, the characteristic radius. We use a Gaussian cut-off radius $r_{\mathrm{t}}=180$ kpc to obtain a finite mass of $M_{\mathrm{h}}=6.3\times 10^{11}\,\mathrm{M_{\odot}}$. The initial exponential stellar disk has the mass $M_{\mathrm{d}}=6.3\times 10^{10}\,\mathrm{M_{\odot}}$ and radial scale-length $R_{0}=2.85$ kpc, and the scale-height $z_{0}=0.6\,\mathrm{kpc}$. Velocities of disk particles are assigned using the epicycle approximation and asymmetric drift correction (Li et al., 2023b). Two models of DM halos are abbreviated as P00 (for $\lambda=0$) and P60 (for $\lambda=0.06$). The angular momentum axis of the DM halo in the P60 model agrees with that of the disk.
@@
Figure 2: Schematic figure of the region where the DM particles are sampled. The figure has the disk edge-on view ($x$-$z$ plane) at $t=0.5$ Gyr. The colored contours are the equipotential surfaces of the total system. The gray dashed box indicates the cylinder with a radius 15 kpc and vertical extension from $z=-15$ to $z=15$ kpc. The dashed ‘circle’ is an equipotential surface with the maximal potential $\Phi_{\mathrm{max}}$ found inside the cylinder. The envelope of the disk is plotted as the solid black curve. We randomly sampled $6.5\times 10^{5}$ DM particles inside $\Phi_{\mathrm{max}}$ equipotential surface whose specific energy is smaller than $\Phi_{\mathrm{max}}$.
@@ 3 Results
Both models, P00 and P60, are subject to the bar instability and their evolution is described in Li et al. (2023b) and in Figure 1, displaying the bar Fourier amplitude, $A_{2}$, the buckling Fourier amplitude, $A_{\mathrm{1z}}$, and the bar pattern speed, $\Omega_{\mathrm{bar}}$. The stellar bar evolution depends on the halo $\lambda$. The P00 bar experiences vertical buckling at $t\sim 3.3$ Gyr, and fully re-grows after, while the P60 bar buckles at about 4 Gyr, essentially leaving behind only an oval distortion, and its torque against the DM, therefore, is very weak.
@@
We have performed an orbit analysis (see e.g., Li et al., 2023c, for details) for both models by constructing a sample of halo particles from each model. The sampled particles have been selected as follows (see Figure 2). (1) We calculated the maximal potential $\Phi_{\mathrm{max}}$ from the region within the cylindrical radius $R=15$ kpc and vertical distance $|z|<15$ kpc at $t=0.5$ Gyr; (2) inside the equipotential surface of $\Phi_{\mathrm{max}}$, we randomly choose $6.5\times 10^{5}$ DM particles whose specific energy is smaller than $\Phi_{\mathrm{max}}$.
@@
We have determined orbital parameters of DM particles, e.g., their maxima along the $z$-axis, $z_{\mathrm{max}}$, and along the cylindrical $R$-axis, $R_{\mathrm{max}}$. By focusing on the evolution of $z_{\mathrm{max}}(t)$, we have detected its decrease by a factor $\geq 2$ over various time. To analyze statistical properties of these orbits, we have binned evolution of $z_{\mathrm{max}}(t)$ into $\Delta t=0.5$ Gyr. Next, we have calculated $\langle z_{\mathrm{max}}\rangle$ — the time-averaged $z_{\mathrm{max}}$ for each consecutive 0.5 Gyr bin from $t=0$ to $t=10$ Gyr. We call the abruptly ‘cooled’ particles those whose $\langle z_{\mathrm{max}}\rangle$ decreases by a factor of $\geq 2$ over $\Delta t\leq 0.5$ Gyr, to separate them from all other cooled halo particles over longer time periods, i.e., 0.5–1.0 Gyr, 1.0–1.5 Gyr, and so on. Note that the $z$-coordinate of each particle has been measured with respect to the Laplace surface defined by the surface where the vertical acceleration $a_{z}$ vanishes (e.g., Dekel & Shlosman, 1983; Li et al., 2023c).
@@
Figure 3: Four representative examples of cooling particles in the P00 model. The $z$-coordinate of each particle is measured with respect to the Laplace surface. $\langle z_{\mathrm{max}}\rangle$ calculated in each 0.5 Gyr bin has been plotted using red curves. (a) Abruptly cooling particle at $\sim 7.5$ Gyr and never heats up again until the end of the run. (b) The particle cools down abruptly at $\sim 3.5$ Gyr and heats up at $\sim 8$ Gyr. (c) The particle experiences multiple coolings at $\sim 2.5$, $\sim 5.8$, and $\sim 8$ Gyr. (d) The particle cools down gradually from $\sim 2$ to $\sim 10$ Gyr.
@@
Figure 4: Statistics of a sub-sample of DM halo particles which cool down from high altitude into the disk plane. These particles are selected under two conditions (see text for details) in the P00 (blue) and P60 (red) models. Top frame: The net number of sub-sampled particles whose $\langle z_{\mathrm{max}}\rangle$ are smaller than the disk thickness after a cooling event. Particles that are heated up and leaving the disk in each time bin are excluded. Bottom frame: The histogram of the duration of individual cooling events for particles that are found in the disk at $t=10$ Gyr (i.e., the particles at the end-point in the top panel).
@@
To obtain the statistics of cooled halo particles, we proceed with the following definitions. First, we assume that the stellar disk extends to $z_{\mathrm{disk}}=3$ kpc above/below the midplane, based on the edge-on disk density contours. Second, we focus on these halo particles which acquire the disk kinematics. Third, in Section 4, we consider a fraction of cooled DM particles as the low metallicity halo stars that have obtained the disk kinematics. For this purpose, we limit our statistical analysis to DM particles with $\langle z_{\mathrm{max}}\rangle\geq 2z_{\mathrm{disk}}$ only before the cooling process begins. This also explains the requirement of cooling $\langle z_{\mathrm{max}}\rangle$ by a factor of $\geq 2$.
We define a sub-sample of all DM halo particles that cool down from the high altitude into the disk in each bin by all types of cooling, i.e., abruptly, gradually, and multiple coolings, $N_{\mathrm{cool}}(t)$. In both the P00 and P60 models, the particles of this sub-sample are filtered out of the total sample discussed above by imposing the two conditions: (1) $\langle z_{\mathrm{max}}\rangle>2z_{\mathrm{disk}}$ at the time of sampling ($t=0.5\>\mathrm{Gyr}$); (2) $\langle z_{\mathrm{max}}\rangle<z_{\mathrm{disk}}$, after the cooling.
The top panel of Figure 4 shows the cumulative number of cooled DM halo particles as a function of time, $N_{\mathrm{cool}}(t)$, i.e., only the DM particles that remain in the disk. The re-heated DM particles, i.e., those leaving the disk, have been excluded. Hence, $N_{\mathrm{cool}}(t=10\,\mathrm{Gyr})$ gives the number of the cooled DM particles remaining in the disk by the end of the run. The bottom panel shows the duration distribution of individual cooling events from particles that are found in the disk at $t=10$ Gyr.
The sample of DM particles used to analyze the statistics of cooling events includes only the particles with $\langle z_{\mathrm{max}}\rangle$ larger than two times $z_{\mathrm{disk}}$ at the time of sampling ($t=0.5$ Gyr). This number is $N_{\mathrm{total}}=325,270$ for the P00 model and 325,019 for the P60 model. To calculate this number, we have determined the total number of sampled DM particles inside the equipotential surface $\Phi_{\mathrm{max}}$. Then corrected it by the fraction of bound particles with $\langle z_{\mathrm{max}}\rangle\geq 2z_{\mathrm{disk}}$.